OpenStudy (anonymous):

at the instant a traffic light turns green, a care starts with a constant acceleration of 6 ft per sec.At the same instant , a truck traveling with a constant velocity of 30 ft per sec passes the car. a) how far beyond its starting point will the car pass the truck? b)how fast will the car be traveling when it passes the truck?

OpenStudy (unklerhaukus):

how long are the vehicles

OpenStudy (anonymous):

it doesnt say:(

OpenStudy (unklerhaukus):

so we have to assume they are point particles/

OpenStudy (anonymous):

yes ..i believe so

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Im not sure

OpenStudy (anonymous):

Acceleration of the car: V(c) = 6t Acceleration of the truck: V(t) = 30 Integrate both to get velocity equations. Integrate both of those to get position equations. Set the position equations equal and solve for t.

OpenStudy (anonymous):

Wait wait. Those are velocity equations. Integrate once.

OpenStudy (anonymous):

so v(c) is 3t^2

OpenStudy (anonymous):

and...v(t) is 30t

OpenStudy (anonymous):

Velocity is the derivative of position, so The integral of a velocity equation is a position equation. You did the integrals correctly, but I would write them as: P(c) = 3t^2 and P(t) = 30t

OpenStudy (anonymous):

Now, there's a middle step here usually. That is, I get a constant of integration. P(c) = 3t^2 + C but C is my starting position, and since you know that your starting position is 0, it drops out. The same thing happens with the truck. If they started at different places, that step would be important.

OpenStudy (anonymous):

Ohh..so if they started at different times, then we would use the C..and the C would just be when they started?

OpenStudy (anonymous):

Right =D

OpenStudy (anonymous):

Or rather, where they started.

OpenStudy (anonymous):

ohhok! ...soo what would we do next?

OpenStudy (anonymous):

Quick example, suppose they tell me that the truck starts 10 feet ahead of the car. I'd get P(T) = 30t + C Well, I know when t = 0, the truck is at position 10, so that gives 30(0) + C = 10 so C = 10 and the position equation is P(T) = 30t + 10

OpenStudy (anonymous):

Well, we have position equations for both vehicles, and we're asked when they pass each other. They pass each other when their positions are the same, so set the position equations equal and solve for t. Let me know what time you get.

OpenStudy (anonymous):

You should get two times, by the way. One of those times just isn't very interesting.

OpenStudy (anonymous):

so is it 3t^2=30t?

OpenStudy (anonymous):

Yes keep going.

OpenStudy (anonymous):

1/10 t^2=t

OpenStudy (anonymous):

what is the next step?

OpenStudy (anonymous):

Been a while since you solved an equation like this?

OpenStudy (anonymous):

It's easier if you gather everything on one side of the equation as your first step: 3t^2 = 30t 3t^2 - 30t = 0

OpenStudy (anonymous):

so its 3t(t-10)=0 so t=0 or t=10?

OpenStudy (anonymous):

Yes =D And as a reminder, any time I have a quadratic equation, that is the highest exponent of the variable is 2, there are two good ways to solve it: factor it out, this way sometimes works, but is often very quick. use the quadratic formula. this way always works, but is a little slower.

OpenStudy (anonymous):

So interpret those results. What does t=0 and t=10 mean? And why would I say that only one of those answers is interesting?

OpenStudy (anonymous):

so, at 10 seconds...both the car and truck are at the same position.And t=10 is interesting..because t=0 is basically at 0 seconds

OpenStudy (anonymous):

Yes. And my question already told me that they pass at that time. I didn't have to do any math to figure that out.

OpenStudy (anonymous):

so the answer to b) is 60 ft/sec

OpenStudy (anonymous):

and the answer to A)300 ft at 10 seconds?

OpenStudy (anonymous):

Yes very good. Use the time you found in the velocity and position equations you wrote.

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

My pleasure =D